OpenStudy (anonymous):

What are the possible values of the four quantum numbers for a 2p electron in boron?

OpenStudy (anonymous):

n is for the shell, l is subshell, ml is the subshell shape, and ms is the spin

OpenStudy (anonymous):

so i'll start it off

OpenStudy (anonymous):

we're only looking at 2p

OpenStudy (anonymous):

so first: n=2, l=1, m=-1, and s=1/2

OpenStudy (anonymous):

It can have a fourth quantum number s that be either +1/2 or -1/2...every shell (n number) has sub-shell (l number) oriented in one of the possible orientations (m number) that only allows two electrons with opposite magnetic spin (s number).

OpenStudy (anonymous):

The neutral boron atom has the following shell configuration:\[[Ne]2s^{2}p^{1}\]2p corresponds to principle quantum number n=2 (2 in 2p), azimuthal quantum number l=1 (p in 2p). I'm almost positive that the choice of magnetic quantum number m and spin quantum number s are arbitrary in this case. To the extent the choice of m and s are arbitrary, any of the following work:\[n=2, l=1, m_{l}=-1, s=+1/2\]\[n=2, l=1, m_{l}=-1, s=-1/2\]\[n=2, l=1, m_{l}=0, s=+1/2\]\[n=2, l=1, m_{l}=0, s=-1/2\]\[n=2, l=1, m_{l}=1, s=+1/2\]\[n=2, l=1, m_{l}=1, s=-1/2\] I recall reading that by convention the lowest available m is assigned first, but +1/2 s electrons are assigned before -1/2 s electrons. I'm having trouble figuring out where I read that. In any case, all of this works out to exactly what BohrMachine posted, above: \[n=2, l=1, m_{l}=-1, s=+1/2\]

OpenStudy (anonymous):

I found my source. Lecture Notes #1 (LN-1), page 9, second paragraph after the bullet points. The Aufbau Principle.